OpenStudy (anonymous):

Verify the equations that appear below. Find the 150thinstance. This means find expressions for functions a(n) and b(n) such that [a(150)]^2-[b(150)]^2= 150^3 and verify the result by showing algebraically that [a(n)]^2-[b(n)]2=n^3 for all n.

OpenStudy (aum):

Doesn't look like this is the complete problem. a(n) and b(n) are not defined. Can you post a screenshot of the question?

OpenStudy (anonymous):

thats whats confusing. I thought the same but that's exactly what the problem says @aum

OpenStudy (anonymous):

sorry i did miss a part! \[1^{2}-0^{2} = 1^3 \] \[3^{2}-1^{2}=2^{3}\]

OpenStudy (anonymous):

@aum

OpenStudy (aum):

\[\large ~~a_n^2 - b_n^2 = n^3 \\ \large ~~1^2 - 0^2 = 1^3 \\ \large ~~3^2 - 1^2 = 2^3 \\ \large ~~6^2 - 3^2 = 3^3 \\ \large 10^2 - 6^2 = 4^3 \\ \]\[\large n:~~ 1,~ 2,~ 3, ~~4,...\\ \large a_n:~ 1,~ 3~, 6, 10, ... ~~~~~~~\normalsize \text {Find the nth term in this sequence.}\\ \large b_n:~~ 0,~ 1,~ 3,~~ 6, ... ~~~~~~~\normalsize \text {Find the nth term in this sequence.} \large \] Find the nth term in the sequence: 1, 3, 6, 10, .... Take the difference in terms: 2 3 4 ... Take the difference again: 1 1 Since the second difference is a constant, this is a quadratic function of the form: \(\large a_n = pn^2 + qn + r\) where p = 1/2 * second difference = 1/2*1=1/2. \(\large a_n = \frac 12n^2 + qn + r\) When \(n = 1, ~a_n = 1\): \(\large 1 = \frac 12*1^2 + q*1 + r = \frac 12 + q + r\) --- (1) When \(n = 2, ~a_n = 3\): \(\large 3 = \frac 12*2^2 + q*2 + r = 2 + 2q + r\) ----- (2) Subtract (1) from (2): \(\large 2 = \frac 32 + q; ~~~~q = \frac 12 \) Put q = 1/2 in (1): \(\large 1 = \frac 12 + \frac 12 + r; ~~~~ r = 0\) Therefore, \(\large a_n = \frac 12n^2 + \frac 12 n = \frac 12 n(n+1)\) Similarly, \(\large b_n = \frac 12n^2 - \frac 12 n = \frac 12 n(n-1)\) Show algebraically, \(\large a_n^2 - b_n^2 = n^3\) \[ \large a_n^2 - b_n^2 = (a_n+b_n)(a_n-b_n) = \\ \large \left( \frac 12n^2 + \frac 12 n + \frac 12n^2 - \frac 12 n \right ) * \left( \frac 12n^2 + \frac 12 n - \frac 12n^2 + \frac 12 n \right ) = \\ \large n^2 * n = n^3 \] Find the 150th instance. That is, set n = 150 and find \(\large a_n\) and \(\large b_n\): \(\large a_{150} = \frac 12 * 150 * 151 = 11,325\) \(\large b_{150} = \frac 12 * 150 * 149 = 11,175\) Make sure \(\large 11325^2 - 11175^2 = 150^3\) \(\large 11325^2 - 11175^2 = 3375000\) \(\large 150^3 = 3375000\)

OpenStudy (anonymous):

wow, i had to go over this a few times for me to understand it. Thanks!